OpenStudy (anonymous):

48x^2+34x-40 factor the trinomial? ( 24+ )( 2- )

OpenStudy (mathstudent55):

First, factor a common factor from all three terms.

OpenStudy (anonymous):

2?

OpenStudy (mathstudent55):

Yes

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

(24x+ )(2x- )?

OpenStudy (mathstudent55):

Then you get 2(24x^2 + 17x - 20)

OpenStudy (mathstudent55):

Now that I wrote the method, let's go over it step by step with your problem.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Your trinomial is 24x^2 + 34x - 20. When compared to ax^2 + bx + c, you see that you have a = 24, b = 17, and c = -20. You follow so far?

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

Ok, Now let's multiply ac, so we get 24*(-20) = -480

OpenStudy (mathstudent55):

Now we need two factors of -480 that add up to 17

OpenStudy (mathstudent55):

How about -32 and 15?

OpenStudy (anonymous):

yess

OpenStudy (mathstudent55):

Sorry, the other way around. -15 and 32 bec they need to add to +17.

OpenStudy (mathstudent55):

-15*32 = -480 -15 + 32 = 17

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Now we factor by parts. That means factor a common factor out of the first two terms, and factor out a common factor out of the last two terms.

OpenStudy (mathstudent55):

We need -15 and 32. Now we rewrite the middle term using these two factors: 24x^2 + 32x - 15x - 20

OpenStudy (mathstudent55):

What do 24x^2 and 32x have in common?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4(6x+8)

OpenStudy (mathstudent55):

4 is the number part. What about x?

OpenStudy (mathstudent55):

You're almost there. 4x(6x + 8)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

One sec. There is a larger number that can be factored.

OpenStudy (anonymous):

8

OpenStudy (mathstudent55):

24x^2 + 32x = 8x(3x + 4)

OpenStudy (mathstudent55):

Good. Now what can be factored out of the last two terms? -15x - 20?

OpenStudy (anonymous):

-5x(3x+4)

OpenStudy (mathstudent55):

Great. Now let's write everything we have so far, includiong the 2 we originally factored out.

OpenStudy (mathstudent55):

BTW, no x this time -15x - 20 = -5(3x + 4)

OpenStudy (mathstudent55):

48x^2 + 34x - 40 = 2(24x^2 + 17x - 20) = 2(24x^2 + 32x - 15x - 20) = 2 [ 8x(3x + 4) - 5(3x + 4) ] The last step now is to pull out the common factor of (3x + 4) inside the brackets: 2 [(3x + 4)(8x - 5) ] = 2(3x + 4)(8x - 5) (Final answer)

OpenStudy (anonymous):

thank you!!

OpenStudy (mathstudent55):

wlcm